At December 31, 2016, Grouper Corporation had the following stock outstanding. 10% cumulative preferred stock, $100 par, 109,101 shares $10,910,100 Common stock, $5 par, 4,021,360 shares 20,106,800 During 2017, Grouper did not issue any additional common stock. The following also occurred during 2017 Income from continuing operations before taxes Discontinued operations (loss before taxes) Preferred dividends declared Common dividends declared Effective tax rate $23,507,100 $3,330,000 $1,091,010 $2,082,400 35 % Compute earnings per share data as it should appear in the 2017 income statement of Grouper Corporation. (Round answers to 2 decimal places, e.g. 1.48.) Earnings Per Share
Answer #1

Answer =

Earning per share
EPS from continuin operation $3.53
EPS from discontinued operation ($0.54)
EPS from net income $2.99

Explanation:

Income from continuing operation before tax $23507100
Income tax (35%) ($8227485)
Income from continuing operation $15279615
Discontinued operations
Loss before taxes $3330000
Less: Applicable income tax (35%) ($1165500)
Loss from discontinued operation ($2164500)
Net income $13115115
Preferred dividend ($1091010)
Income available for common shareholders $12024105
Number of shares of common stock outstanding 4021360

1)EPS from continuing operation = ($15279615-$1091010)/4021360 = $3.53

2) EPS from discontinued operation = -$2164500/4021360 = ($0.54)

3) EPS from net income = $12024105/4021360 = $2.99

I have given complete and detailed solution to your problem, in case of any query you can comment and

PLEASE DO GIVE POSITIVE RATING